Three point charges are arranged along the y-axis in a vacuum. The topmost charge bears a charge of -4.0 , the middle charge has a charge of 3.0 , and the bottom carries a charge of -7.0 . What is the magnitude and direction of the net electrostatic force that the middle charge experiences if top-middle is separated by 0.2 m of distance and middle-bottom is also separated by 0.2 m of distance?

1
Expert's answer
2022-02-06T14:33:58-0500

"F=F_1-F_2=\frac{kq}{r^2}(q_1-q_2)=2.03~N."